# Causal Discovery Inspired Unsupervised Domain Adaptation for Emotion-Cause Pair Extraction

### v1

@@ Abstract
This paper tackles the task of emotion-cause pair extraction in the unsupervised domain adaptation setting. The problem is challenging as the distributions of the events causing emotions in target domains are dramatically different than those in source domains, despite the distributions of emotional expressions between domains are overlapped. Inspired by causal discovery, we propose a novel deep latent model in the variational autoencoder (VAE) framework, which not only captures the underlying latent structures of data but also utilizes the easily transferable knowledge of emotions as the bridge to link the distributions of events in different domains. To facilitate knowledge transfer across domains, we also propose a novel variational posterior regularization technique to disentangle the latent representations of emotions from those of events in order to mitigate the damage caused by the spurious correlations related to the events in source domains. Through extensive experiments, we demonstrate that our model outperforms the strongest baseline by approximately 11.05% on a Chinese benchmark and 2.45% on a English benchmark in terms of weighted-average F1 score. The source code will be publicly available upon acceptance.
@@ 1 Introduction
It is a common practice to project texts into latent representations for improving language understanding Wang et al. (2019). Existing techniques disentangle different types of latent representations by applying regularization terms to enforce independence between the corresponding random variables Cheng et al. (2020). However, the independence assumption contradicts the fact that emotions and the events causing them are statically dependent.
@@ 2 Challenges in UDA-ECPE
To provide an intuitive understanding of the above mentioned challenges in the UDA setting, we visualize the clause embeddings, namely $p({\mathbf{Z}}^{c})$, for ground-truth emotion and emotion causes respectively on CH-ECPE and EN-ECPE, and compare them with the sentence embeddings for a widely used domain adaptation corpus Amazon Reviews Blitzer et al. (2007) using t-SNE. As the original CH-ECPE are not partitioned based on domains, we manually assign each data point in the corpus with the corresponding domain label. Further details are provided in Sec. 4.1.
As shown in Figure 5, the data points of Chinese emotion clauses from various CH-ECPE’s domains are strongly overlapped, the domain divergences are far smaller than those of the embeddings of the emotion causes. It is thus challenging for existing UDA methods, which work only in the cases that the distribution shift from a source domain to a target domain is small, as illustrated in Fig.2(a) Zhao et al. (2019); Kumar et al. (2020). In addition, we employ two different datasets as different domains for English. For English corpora similar tendency can be found in A.1.
@@
Figure 2: The t-SNE visualizations of the sentence embeddings from Amazon Reviews multi-domain sentiment corpus, the clause embeddings from the Chinese UDA-ECPE corpora for English UDA-ECPE corpora please refer to A.1
@@ 3 Methodology
In theory, the corresponding divergence measures $\mathbb{D}_{k}(q({\mathbf{Z}}^{e}|{\mathbf{X}}_{ij})\|q({\mathbf{Z}}^{c}|{\mathbf{X}}_{ij}))$ should not assume absolute continuity Royden and Fitzpatrick (1988), which requires that $q(Z^{e}_{i}|{\mathbf{X}}_{ij})>0$ for every $q(Z^{c}_{i}|{\mathbf{X}}_{ij})>0$, vice versa. In reality, a random variable $Z_{i}^{e}$ may have high probability in the region where a $Z_{j}^{c}$ has zero probability. To tackle this, we choose Bhattacharyya distance Bhattacharyya (1946) and maximum mean discrepancy (MMD) Gretton et al. (2012) respectively as a regularizer. Each of them has its own strength. More details are covered in Sec. 3.2.
@@ 3.2.1 Source Domain Training
where $\phi$ is a mapping function that projects both ${\bm{z}}^{e}$ and ${\bm{z}}^{c}$ into a reproducing kernel Hilbert space denoted by $\mathcal{H}$. In this work, we mainly adopt this regularizer in experiments due to its superior performance over the other two.
@@ 4.1 Experimental Setup
Datasets. Since there is no corpus for ECPE in the UDA setting, we divide CH-ECPE into multiple domains. Given the fact that the documents in CH-ECPE are Chinese news articles sampled from the THUCNews dataset Li and Sun (2007), we employ the topic classifier THUCTC Sun et al. (2016) trained on the THUCNews dataset to categorize CH-ECPE into 14 subsets based on topics and choose the largest five as the final domains (e.g. home, society and finance, etc.). To further improve the purity of classification, based on THUCTC’s classification results, we conduct manual inspection and labeling to complete the domain classification of CH-ECPE. Also, in the English language setting, we view EN-ECPE and Recognizing Emotion Cause in CONversations (RECCON) Poria et al. (2021) – an English dataset specifically designed for identifying the causes of emotions within conversations, as the two source-target domains. Table 4 summarizes the statistics of each corpus and can be found in A.2.
@@
Baselines. To make a fair comparison, we adapt the three existing ECPE models RankCP, UTOS, UECA-Prompt (all employ BERT as the backbone model) for emotion extraction (EE) and ECPE. In addition, since the universal prompt-based method for ECA tasks (UECA-Prompt) Zheng et al. (2022) is designed to solve the different Emotion cause analysis (ECA) tasks in an unified framework, we thus only integrate three UDA approaches on the two ECPE models (RankCP Wei et al. (2020) and UTOS Cheng et al. (2021)) in the ECPE task to further demonstrate the effectiveness of our model. The introduction of baseline method and implementation detail please refer to A.2.
@@ 4.2 Results and Analysis
  Model Society $\rightarrow$ Home Society $\rightarrow$ Finance Society $\rightarrow$ Education Society $\rightarrow$ Entertainment Weighted Average EE (%) ECPE (%) EE (%) ECPE (%) EE (%) ECPE (%) EE (%) ECPE (%) EE (%) ECPE (%) $P$ $R$ $F1$ $P$ $R$ $F1$ $P$ $R$ $F1$ $P$ $R$ $F1$ $P$ $R$ $F1$ $P$ $R$ $F1$ $P$ $R$ $F1$ $P$ $R$ $F1$ $F1$ $F1$ (a) S: Society RankCP 21.90 25.22 23.44 13.14 14.54 13.80 18.04 21.00 19.41 8.56 9.86 9.17 26.13 31.90 28.73 18.59 22.29 20.27 26.87 32.73 29.51 13.43 16.36 14.75 23.49 13.65 RankCP+Ada-TSA 18.55 21.16 19.77 12.30 13.48 12.86 15.86 17.44 16.61 7.12 7.75 7.42 20.62 24.54 22.41 11.86 13.86 12.78 23.44 27.27 25.21 6.25 7.27 6.72 19.65 11.41 RankCP+DANN 91.51 98.15 94.72 51.69 93.85 66.67 85.06 93.24 88.96 40.38 75.35 52.58 82.87 92.02 87.21 43.01 74.10 54.42 77.78 89.09 83.05 30.48 58.18 40.00 92.03 60.93 RankCP+MEDM 20.17 23.12 21.55 12.77 14.07 13.39 20.43 23.84 22.00 9.76 11.27 10.46 24.14 30.06 26.78 13.30 16.27 14.63 14.52 16.36 15.38 6.45 7.27 6.84 22.04 12.63 UTOS 91.51 47.72 62.73 70.99 35.58 47.40 93.33 49.82 64.97 71.33 37.68 49.31 92.21 43.56 59.17 67.09 31.93 43.27 71.43 27.27 39.47 47.62 18.18 26.32 61.77 46.39 UTOS+Ada-TSA 18.55 21.16 19.77 12.30 13.48 12.86 15.86 17.44 16.61 7.12 7.75 7.42 20.62 24.54 22.41 11.86 13.86 12.78 23.44 27.27 25.21 6.25 7.27 6.72 19.65 11.41 UTOS+DANN 84.96 61.13 71.10 56.41 40.07 46.86 89.55 64.06 74.69 57.84 41.55 48.36 86.92 57.06 68.89 62.28 42.77 50.71 80.65 45.45 58.14 48.39 27.27 34.88 71.04 47.16 UTOS+MEDM 52.80 55.60 54.16 14.63 33.32 20.31 15.31 89.68 26.15 0.64 13.03 1.21 53.00 62.50 46.01 24.23 28.31 26.11 57.50 41.82 48.42 12.64 20.00 15.49 46.82 16.70 UECA-Prompt 75.59 74.66 75.12 50.92 61.43 55.69 71.01 69.75 70.38 51.13 62.63 56.30 75.84 82.82 79.17 48.84 62.87 54.97 73.58 70.91 72.22 45.21 60.00 51.56 74.48 55.55 Ours 81.77 76.14 78.85 58.59 71.98 64.60 86.42 81.49 83.88 75.96 82.01 78.87 83.85 82.82 83.33 74.30 79.64 76.88 86.00 78.18 81.90 84.62 80.00 82.24 80.63 71.35 (b) S: Home Home $\rightarrow$ Society Home $\rightarrow$ Finance Home $\rightarrow$ Education Home $\rightarrow$ Entertainment RankCP 83.88 91.82 87.67 44.33 75.42 55.84 86.56 93.95 90.10 43.41 75.35 55.08 83.33 92.02 87.46 44.48 77.71 56.58 84.48 89.09 86.73 36.78 58.18 45.07 81.85 51.31 RankCP+Ada-TSA 16.38 19.37 17.75 8.25 9.51 8.84 20.42 24.20 22.15 8.11 9.51 8.75 18.82 21.47 20.06 10.75 12.05 11.36 22.06 27.27 24.39 5.88 7.27 6.50 18.01 8.40 RankCP+DANN 29.29 37.45 32.87 26.79 33.43 29.74 25.00 29.54 27.08 14.76 17.25 15.91 31.34 41.72 35.79 17.51 22.89 19.84 27.27 32.73 29.75 13.64 16.36 14.88 29.49 22.73 RankCP+MEDM 15.43 17.36 16.34 6.89 7.55 7.20 7.61 7.47 7.54 2.17 2.11 2.14 22.04 25.15 23.50 8.60 9.64 9.09 23.81 27.27 25.42 6.35 7.27 6.78 14.55 5.81 UTOS 88.56 51.08 64.79 70.69 40.08 51.16 90.00 57.65 70.28 62.30 40.14 48.82 92.13 50.31 65.08 70.79 37.95 49.41 78.26 32.73 46.15 52.17 21.82 30.77 60.57 45.89 UTOS+Ada-TSA 16.38 19.37 17.75 8.25 9.51 8.84 20.42 24.20 22.15 8.11 9.51 8.75 18.82 21.47 20.06 10.75 12.05 11.36 22.06 27.27 24.39 5.88 7.27 6.50 18.01 8.40 UTOS+DANN 87.98 62.98 73.41 63.04 44.62 52.25 89.36 59.79 71.64 63.16 42.25 50.63 85.32 57.06 68.38 60.91 40.36 48.55 78.12 45.45 57.47 37.50 21.82 27.59 66.45 46.63 UTOS+MEDM 33.96 65.28 44.67 5.52 37.20 9.61 13.85 92.88 24.11 0.61 14.44 1.16 39.21 54.60 45.64 6.45 30.12 10.63 46.67 50.91 48.70 9.3 21.82 13.04 37.31 7.37 UECA-Prompt 76.52 85.08 80.57 66.33 63.11 64.68 78.04 82.21 80.07 61.96 59.17 60.53 75.14 81.60 78.24 66.27 65.87 66.07 75.93 74.55 75.23 58.18 58.18 58.18 69.10 59.04 Ours 86.07 79.77 82.80 68.78 75.07 71.79 81.79 84.70 83.22 76.03 83.39 79.54 80.72 82.21 81.46 84.71 79.64 82.10 84.31 78.18 81.13 83.33 81.82 82.57 76.72 70.09
@@
Table 2: Experimental results of our models and the baseline models on EN-ECPE and RECCON.
@@
Overall Comparisons.
Table 1 and Table 2 report the results of our models and the baselines on the ECPE task, as well as the EE subtask. To dispel the doubt that our model outperforms the baselines only because they are developed in the supervised setting, we apply the SOTA UDA methods Ada-TS Zhang et al. (2021), DANN Ganin et al. (2016) and MEDM Wu et al. (2021) to the two baselines RankCP and UTOS on the UDA-ECPE task. MEDM is a minimal-entropy UDA approach that introduces diversity maximization to regulate entropy minimization for seeking a close-to-ideal domain adaptation. Ada-TSA is a recently proposed adapter-based UDA approach in which the newly-added adapters can capture transferable features between source and target domains by using the domain-fusion scheme. DANN is a widely adopted adversarial-based UDA approach that learns domain invariant representations through a domain discriminator. It can be found that after applying the UDA framework, RankCP and UTOS significantly improved their performance and became comparable with the SOTA prompt-based model UECA-Prompt.
However, though we employ UDA (for RankCP and UTOS) while leverage the powerful ability of the Large Language Model (LLM) (for UECA-Prompt) to enhance the baseline models, the baseline models still perform worse than our proposed model. On CH-ECPE, our model outperforms the RankCP+DANN by $10.42\%$ when treating society as the source domain, and UECA-Prompt by $11.05\%$ with home as the source domain in terms of weighted average F1. On EN-ECPE, our model is better than the supervised learning model RankCP by $2.45\%$. Also, we can observe that our models get the best ECPE results in almost all of the domains except the $Society\rightarrow Home$ setting, indicating the generalization ability of the proposed approach. It is worth mentioning that our model performs the best even it does not always achieve the best performance on the EE subtask. Note that there is a significant performance gap between the Chinese and English benchmarks. The cause of this gap mainly due to the distribution bias problem where the five domains used for testing in the Chinese benchmark are extracted from the same corpus, i.e., CH-ECPE, however the two domains under the English setting derive from the two different datasets RECCON and EN-ECPE. Therefore, compared with the Chinese domains, the two English domains share less knowledge between each other, making the model hard to transfer from one domain to another. Overall, the results demonstrate the strengths of our model in terms of identifying new causal relations between events and emotions in new domains.
@@
Also, we remove Emotion and Event adapters and use the unified pair representation as the input for both the emotion and event encoders. By doing this we lost performance for all domains, as the Table 3 shows. It is proved that using the different vectors to represent the emotion / event variables is a better solution. In addition, we also conduct experiments on investigating the efficacy of self-training and regularizer, detailed in A.3.

### v2

@@ Abstract
This paper tackles the task of emotion-cause pair extraction in the unsupervised domain adaptation setting. The problem is challenging as the distributions of the events causing emotions in target domains are dramatically different than those in source domains, despite the distributions of emotional expressions between domains are overlapped. Inspired by causal discovery, we propose a novel deep latent model in the variational autoencoder (VAE) framework, which not only captures the underlying latent structures of data but also utilizes the easily transferable knowledge of emotions as the bridge to link the distributions of events in different domains. To facilitate knowledge transfer across domains, we also propose a novel variational posterior regularization technique to disentangle the latent representations of emotions from those of events in order to mitigate the damage caused by the spurious correlations related to the events in source domains. Through extensive experiments, we demonstrate that our model outperforms the strongest baseline by approximately 11.05% on a Chinese benchmark and 2.45% on a English benchmark in terms of weighted-average F1 score. We have released our source code and the generated dataset publicly at: https://github.com/tk1363704/CAREL-VAE.
@@ 1 Introduction
It is a common practice to project texts into latent representations for improving language understanding Wang et al. (2019). Existing techniques disentangle different types of latent representations by applying regularization terms to enforce independence between the corresponding random variables Cheng et al. (2020). However, the independence assumption contradicts the fact that emotions and the events causing them are statically dependent. Furthermore, Large Language Models (LLMs), as a powerful general-purpose natural language processing tool, should be well-suited for the UDA-ECPE task. However, in many related studies Gao et al. (2019); Zecevic et al. (2023); Kıcıman et al. (2023); Romanou et al. (2023); Jacovi et al. (2023); Gao et al. (2023); Jin et al. (2024); Feng et al. (2024), researchers have found that LLMs are not particularly effective at solving causal discovery tasks.
@@ 2 Challenges in UDA-ECPE
To provide an intuitive understanding of the above challenges in the UDA setting, we visualize the clause embeddings, namely $p({\mathbf{Z}}^{c})$, for ground-truth emotion and emotion causes respectively on CH-ECPE and EN-ECPE, and compare them with the sentence embeddings for a widely used domain adaptation corpus Amazon Reviews Blitzer et al. (2007) using t-SNE. As the original CH-ECPE are not partitioned based on domains, we manually assign each data point in the corpus with the corresponding domain label. Further details are provided in Sec. 4.1.
As shown in Figure 2, the data points of Chinese emotion clauses from various CH-ECPE’s domains are strongly overlapped, the domain divergences are far smaller than those of the embeddings of the emotion causes. It is thus challenging for existing UDA methods, which work only in the cases that the distribution shift from a source domain to a target domain is small, as illustrated in Fig.2(a) Zhao et al. (2019); Kumar et al. (2020).
In addition, we employ two different datasets as different domains for English. The English corpora similar tendency can be found in A.1. As shown in Fig.4, regardless if a clause mentions an emotion or an emotion cause, there is a very clear boundary between the two domains. Their domain differences are largely caused by the differences between the two datasets.
@@
Figure 2: The t-SNE visualizations of the sentence embeddings from Amazon Reviews multi-domain sentiment corpus and the clause embeddings from the Chinese UDA-ECPE corpora.
@@ 3 Methodology
In theory, the corresponding divergence measures $\mathbb{D}_{\mathrm{KL}}(q({\mathbf{Z}}^{e}|{\mathbf{X}}_{ij})\|q({\mathbf{Z}}^{c}|{\mathbf{X}}_{ij}))$ should not assume absolute continuity Royden and Fitzpatrick (1988), which requires that $q(Z^{e}_{i}|{\mathbf{X}}_{ij})>0$ for every $q(Z^{c}_{i}|{\mathbf{X}}_{ij})>0$, vice versa. In reality, a random variable $Z_{i}^{e}$ may have high probability in the region where a $Z_{j}^{c}$ has zero probability. To tackle this, we choose Bhattacharyya distance Bhattacharyya (1946) and maximum mean discrepancy (MMD) Gretton et al. (2012) respectively as a regularizer. Each of them has its own strength. More details are covered in Sec. 3.2.
@@ 3.2.1 Source Domain Training
where $\phi$ is a mapping function that projects both ${\bm{z}}^{e}$ and ${\bm{z}}^{c}$ into a reproducing kernel Hilbert space denoted by $\mathcal{H}$. In this work, we mainly adopt this regularizer in experiments due to its superior performance over the other two. More in-depth discussion about the design of $Omega^{\text{MMD}}$ can be found in Appendix A.1.
@@ 4.1 Experimental Setup
Datasets. Since there is no corpus for ECPE in the UDA setting, we divide CH-ECPE into multiple domains. Given the fact that the documents in CH-ECPE are Chinese news articles sampled from the THUCNews dataset Li and Sun (2007), we employ the topic classifier THUCTC Sun et al. (2016) trained on the THUCNews dataset to categorize CH-ECPE into 14 subsets based on topics and choose the largest five as the final domains (e.g. home, society and finance, etc.). To further improve the purity of classification, based on THUCTC’s classification results, we conduct manual inspection and labeling to complete the domain classification of CH-ECPE. Also, in the English language setting, we view EN-ECPE and Recognizing Emotion Cause in CONversations (RECCON) Poria et al. (2021) – an English dataset specifically designed for identifying the causes of emotions within conversations, as the two source-target domains. Table 4 (in A.3) summarizes the statistics of each corpus.
@@
Baselines.
To make a fair comparison, we adapt the three existing ECPE models RankCP, UTOS, UECA-Prompt (all employ BERT as the backbone model) for emotion extraction (EE) and ECPE. In addition, since the universal prompt-based method for ECA tasks (UECA-Prompt) Zheng et al. (2022) is designed to solve the different Emotion cause analysis (ECA) tasks in an unified framework, we thus only integrate three UDA approaches on the two ECPE models (RankCP Wei et al. (2020) and UTOS Cheng et al. (2021)) in the ECPE task to further demonstrate the effectiveness of our model. The introduction of baseline method and implementation detail please refer to A.3.
@@ 4.2 Results and Analysis
  Model EE (%) ECPE (%) EE (%) ECPE (%) EE (%) ECPE (%) EE (%) ECPE (%) EE (%) ECPE (%) $F1$ $F1$ $F1$ $F1$ $F1$ $F1$ $F1$ $F1$ $F1$ $F1$ (a) S: Society Society $\rightarrow$ Home Society $\rightarrow$ Finance Society $\rightarrow$ Education Society $\rightarrow$ Entertainment Weighted Average RankCP 23.44 13.80 19.41 9.17 28.73 20.27 29.51 14.75 23.49 13.65 RankCP+Ada-TSA 19.77 12.86 16.61 7.42 22.41 12.78 25.21 6.72 19.65 11.41 RankCP+DANN 94.72 66.67 88.96 52.58 87.21 54.42 83.05 40.00 92.03 60.93 RankCP+MEDM 21.55 13.39 22.00 10.46 26.78 14.63 15.38 6.84 22.04 12.63 UTOS 62.73 47.40 64.97 49.31 59.17 43.27 39.47 26.32 61.77 46.39 UTOS+Ada-TSA 19.77 12.86 16.61 7.42 22.41 12.78 25.21 6.72 19.65 11.41 UTOS+DANN 71.10 46.86 74.69 48.36 68.89 50.71 58.14 34.88 71.04 47.16 UTOS+MEDM 54.16 20.31 26.15 1.21 46.01 26.11 48.42 15.49 46.82 16.70 UECA-Prompt 75.12 55.69 70.38 56.30 79.17 54.97 72.22 51.56 74.48 55.55 Ours 78.85 64.60 83.88 78.87 83.33 76.88 81.90 82.24 80.63 71.35 (b) S: Home Home $\rightarrow$ Society Home $\rightarrow$ Finance Home $\rightarrow$ Education Home $\rightarrow$ Entertainment Weighted Average RankCP 87.67 55.84 90.10 55.08 87.46 56.58 86.73 45.07 81.85 51.31 RankCP+Ada-TSA 17.75 8.84 22.15 8.75 20.06 11.36 24.39 6.50 18.01 8.40 RankCP+DANN 32.87 29.74 27.08 15.91 35.79 19.84 29.75 14.88 29.49 22.73 RankCP+MEDM 16.34 7.20 7.54 2.14 23.50 9.09 25.42 6.78 14.55 5.81 UTOS 64.79 51.16 70.28 48.82 65.08 49.41 46.15 30.77 60.57 45.89 UTOS+Ada-TSA 17.75 8.84 22.15 8.75 20.06 11.36 24.39 6.50 18.01 8.40 UTOS+DANN 73.41 52.25 71.64 50.63 68.38 48.55 57.47 27.59 66.45 46.63 UTOS+MEDM 44.67 9.61 24.11 1.16 45.64 10.63 48.70 13.04 37.31 7.37 UECA-Prompt 80.57 64.68 80.07 60.53 78.24 66.07 75.23 58.18 69.10 59.04 Ours 82.80 71.79 83.22 79.54 81.46 82.10 81.13 82.57 76.72 70.09
Table 1: Experimental results of our models and baselines utilizing F1 score (F1) as metrics on the ECPE and EE tasks on the Chinese dataset. Emotion Extraction is denoted by EE while ECPE refers to Emotion-Cause Pair Extraction. S refers to source domain.
@@
Table 2: Experimental results of our models and the baseline models on the English ECPE and EE tasks.
Overall Comparisons.
Table 1 (the ECPE and EE tasks on the Chinese dataset) and Table 2 (the ECPE and EE tasks on the English dataset) report the results of our models and the baselines on the ECPE task, as well as the EE subtask. For more detailed metrics on , i.e., precision (P) and recall (R) on Chinese dataset, please refer to Table 5 in the Appendix section. In Table 1 and 2, the “Weighted Average ECPE” metric is the most important indicator of the model’s cross-domain ECPE capability.
To dispel the doubt that our model outperforms the baselines only because they are developed in the supervised setting, we apply the SOTA UDA methods Ada-TS Zhang et al. (2021), DANN Ganin et al. (2016) and MEDM Wu et al. (2021) to the two baselines RankCP and UTOS on the UDA-ECPE task. MEDM is a minimal-entropy UDA approach that introduces diversity maximization to regulate entropy minimization for seeking a close-to-ideal domain adaptation. Ada-TSA is a recently proposed adapter-based UDA approach in which the newly-added adapters can capture transferable features between source and target domains by using the domain-fusion scheme. DANN is a widely adopted adversarial-based UDA approach that learns domain invariant representations through a domain discriminator. It can be found that after applying the UDA framework, RankCP and UTOS significantly improved their performance and became comparable with the SOTA prompt-based model UECA-Prompt.
However, though we employ UDA (for RankCP and UTOS) while leverage the powerful ability of the Large Language Model (LLM) (for UECA-Prompt) to enhance the baseline models, the baseline models still perform worse than our proposed model. On CH-ECPE, our model outperforms the RankCP+DANN by $10.42\%$ when treating society as the source domain, and UECA-Prompt by $11.05\%$ with home as the source domain in terms of weighted average F1. On EN-ECPE, our model is better than the supervised learning model RankCP by $2.45\%$. Also, we can observe that our models get the best ECPE results in almost all of the domains except the $Society\rightarrow Home$ setting, indicating the generalization ability of the proposed approach. It is worth mentioning that our model performs the best even it does not always achieve the best performance on the EE subtask.
Note that there is a significant performance gap between the Chinese and English benchmarks. The cause of this gap mainly due to the distribution bias problem where the five domains used for testing in the Chinese benchmark are extracted from the same corpus, i.e., CH-ECPE, however the two domains under the English setting derive from the two different datasets RECCON and EN-ECPE. Therefore, compared with the Chinese domains, the two English domains share less knowledge between each other, making the model hard to transfer from one domain to another.
To sum up, for the Weighted Average ECPE metric, which is the most important indicator of the model’s cross-domain ECPE capability, RankCP and UTOS, even when using domain adaptation techniques (MEDM, ada-TSA, and DANN), performed lower than our model on both Chinese and English datasets. This demonstrates that our method can effectively enable the model to learn cross-domain ECPE capabilities, and prove the strengths of our model in terms of identifying new causal relations between events and emotions in new domains.
@@
Also, we remove Emotion and Event adapters and use the unified pair representation as the input for both the emotion and event encoders. By doing this we lost performance for all domains, as the Table 3 shows. It is proved that using the different vectors to represent the emotion / event variables is a better solution. In addition, we also conduct experiments on investigating the efficacy of self-training and regularizer, detailed in A.4.